COLOR-CENTER LASER SPECTRUM OF FLUOROBENZENE OBSERVED WITH THE PULSED SUPERSONIC JET EXPANSION TECHNIQUE

\underline{M. ISHIGURO}, K. HARADA, K. TANAKA, and T. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Higashiku, Fukuoka 812-8581, Japan..

The infrared spectrum of fluorobenzene in the CH-stretch region (3040-3120 cm^{−1}) has been observed with a color-center laser spectrometer combined with a pulsed supersonic jet expandion technique. Due to the rotational as well as cooling in the jet, well resolved rovibrational lines were recorded. At least 5 fundamental bands (a_{1}; ν_{2}, ν_{13}, ν_{20a} b_{1}; ν_{7b}, ν_{20b}) are expected in this region according to the opto-thermal Raman conversion dye laser spectroscopy ^{a}. So far two parallel bands (ν_{13}, ν_{20a}) were analyzed to confirm the vibrational assignments given by the low-resolution spectroscopy. The rotational constants and band origins were derived from the observed spectra assuming those for ground state to the microwave results ^{b}^{,} ^{c}. The ν_{20a} band were found to be overlapped additionally with the two vibrational bands presumably caused by higher order Fermi interaction, whose inertia defects are small negative about −0.5 amuÅ^{2}. Their vibrational assignments are obscure yet, but the negative inatria defects imply that they include two or four quanta of low frequency out-of-plane vibrational modes such as ν_{11}(ω = 248 cm^{−1}) and ν_{16b}(498 cm^{−1}).

aM. Scotoni, S. Oss, L. Lubich, S. Furlani, and D. Bassi, J. Chem. Phys. \underline{103}, 897 (1995)

bM. Onda, H. Mukaida, and I. Yamaguchi, J. Mol. Spectrosc. \underline{176}, 146 (1996)

cL. Nygaard, I. Bojesen, T. Pedersen, and J. R. Andersen. J. Mol. Struct. \underline{2}, 209 (1968)